HUBBLE SPACE TELESCOPE DAILY REPORT #1013 PERIOD COVERED: 0800 EST Daily Status Report as of 315/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) The Planetary Camera was used to make standard Earth calibrations images for flat fielding of PC observations. Medium and narrow filters were used, with and without neutral density filters. Four observations were executed as scheduled. Two observations were reported underexposed, and two were reported to be normal. 1.2 Completed Two Sets of PC 4929 (WF/PC Cycle 2 :Improved Dark Calibration) The Planetary Camera was used to make two dark calibration images. The two exposures lasted 1200 seconds each, and the camera was pre-flashed before the observations. Both observations were reported to be normal exposures. 1.3 Completed HSP/VIS 4569 (A Search for Microflaring Activity in Dwarf M Stars _ Cycle 3 High)) The High Speed Photometer was used to make observations of variable dwarf star YZ-Cmi to look for microflares which has a short duration and occurs with numbers inversely proportional to the flare energy. A total of five images were received and displayed. They were reported to be satisfactory. There were three short recentering events reported during the observations. 1.4 Completed Three Sets of WFC 5013 (WF/PC Cycle 3 Calibration: WFC Flats Narrow. 2) The Wide Field Camera was used to make Earth flat calibration images to be used for flat fielding of WFC observations. The observations used the narrow band filters, both with and without neutral density filters. A total of nine observations were received. One image was saturated, two were underexposed, four were reported to be okay, and one still remains on the science tape recorder. 1.5 Completed WFC 4381 (HST Medium - Deep Survey - Cycle 3) The Wide Field Camera was used to make three sets of medium-deep high latitude sky survey observations taken in parallel with the prime science observations. The observations were executed as scheduled, and no problems were reported. 1.6 Completed WFC 5075 (WF/PC Cycle 3 Calibration: WFC Flats Wide. 1) Observations were made to make a standard calibration to obtain Earth flats using the Wide Field Camera. A total of four observations were taken. Four observations were received and displayed. One image was saturated, one was underexposed, and two were reported to be okay. 1.7 Completed GHRS/2 3472 (The Properties of Single Interstellar Clouds - Cycle 2) The High Resolution Spectrograph was used to make Echelle spectral observations of the broad-lined, early-type star HD037128. Both the Echelle and grating G160M were used. A total of six observations were executed, received and displayed without problems. 1.8 Completed FOS/RD 4818 (Quasar Absorption Line Survey - Cycle 2: Repeat for HOPR#0089) The Faint Object spectrograph (Red side) was used to make observations of quasar HS0624+6907 to obtain the profiles of its spectral lines. This covered from 1600 A0 to 3300 A0. The observations were monitored in the OSS, and the data quality was assessed. No anomalies were reported. 1.9 Completed HSP/UV2 3255 (Search for Optical Variability Associated With Black Holes) The High speed Photometer was used to make ultraviolet observations of the X-Ray binary star A0620-00 who's companion is expected to be a black hole. Observations were made to look for variations in the millisecond time scale caused by matter spiraling toward the event horizon in the accretion disk surrounding the black hole. The observations were executed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 5 Scheduled Reacquisitions : 12 Successful : 11 Primary Guide Star Acquisition failed at 315/0542Z during a ZOE. The subsequent reacquisition also failed during the ZOE. The follow acquisition was successful. 2.2 FHST Updates: Scheduled : 0 Successful : 0 2.3 Significant HSTARs: None 2.4 Operations Note: Simulation JIS #6 (EVA 4 and 5, plus Deploy) was initiated Monday morning, Nov. 8, 1993 at 8 AM and was scheduled to last for 59 hours. It was successfully terminate yesterday, Wednesday, Nov. 11, around 8 PM. All test objective were met. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete WFC 5076 (WF/PC Cycle 3 Calibration: WFC Flats Wide. 2) 3.2 Complete WFC 5013 (WF/PC cycle 3 Calibration: WFC Flats Narrow. 2) 3.3 Complete PC 5072 (WF/PC Cycle 3 Calibration: PC Flats Wide. 1) 3.4 Complete FOS/RD and GHRS/2 4685 (Spectra and Chromospheres of Carbon and M Type Stars - Cycle 3 High) 3.5 Complete HSP/POL 4889 (Gravitational Leneses Part I - Augmentation - Cycle 3 High) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-